OpenStudy (anonymous):

if y1=cos4x is a solution to the DE: y''+16y=0. Find y2

OpenStudy (anonymous):

Find a y2 that is also a solution

OpenStudy (anonymous):

\[y_2=y_1\int{}{}\frac{e^\int{}{}{p(x)dx}}{y_1^2}\] or use substitution y=u(x)y1

OpenStudy (anonymous):

y=ucos4x y'=-4usin4x+cos4xu'

OpenStudy (anonymous):

it's supposed to reduce to a order 1 linear: y'+p(x)y=0 using the y=ux

OpenStudy (anonymous):

y=uy1 i mean

OpenStudy (anonymous):

the integral way you get a crazy integral

OpenStudy (anonymous):

wait i might know what happened

OpenStudy (anonymous):

ahh i wrote it down wrong on my paper but here it is right

OpenStudy (anonymous):

y''=-16ucos4x-4u'sin4x+cos4xu''-4u'sin4x=-16ucos4x-8u'sin4x+cos4xu''

OpenStudy (anonymous):

y''+16y=-16ucos4x-8u'sin4x+cos4xu''+16ucos4x=-8u'sin4x+cos4xu''

OpenStudy (anonymous):

let w=u' cos(4x)w'-8sin(4x)w=0

OpenStudy (anonymous):

now it's in linear

OpenStudy (anonymous):

\[w'\frac{-8sin(4x)}{cos(4x)}w=0\] \[w'-8tan(4x)w=0\]

OpenStudy (anonymous):

y2=sin(4x)

OpenStudy (anonymous):

how'd you get it

OpenStudy (anonymous):

\[e^{\int{}{}-8tan(4x)}\]

OpenStudy (anonymous):

\[e^{\int{}{}-2tan(u)du}\]= \[e^{-2(-ln|cos u|)}\]= \[e^{ln|cosu^2|}\]= \[cos^2u=cos^2(4x)\]

OpenStudy (anonymous):

Just replace y by sin(4x) and y'' by -16 sin(4x)

OpenStudy (anonymous):

that's not showing anyy work at all